# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ A.1 Notation, statistical model and assumptions
$\overline{f}({\bm{x}},{\bm{g}})=\mathbb{E}_{\mathbf{u}}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})].$ (24)
@@
It will be useful later when talking about the posterior mean of $\overline{f}$ to be able to freely interchange the order of integration over $f$ and ${\mathbf{u}}$. This is possible by Fubini’s Theorem / the law of total expectation, provided that the expectation $\mathbb{E}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})]$ over both jointly is well defined.
@@
We can exchange the order of integration over $f$ and ${\mathbf{u}}$ when calculating the posterior mean of $\overline{f}$. That is,
$\overline{\mu}^{n}({\bm{x}},{\bm{g}})=\mathbb{E}\bigl{[}\,\overline{f}({\bm{x}},{\bm{g}})\,\big{|}\,\mathcal{F}^{n}\bigr{]}=\mathbb{E}_{\mathbf{u}}\bigl{[}\mu^{n}({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})\bigr{]}.$
@@
The sample paths of $f$ are continuous and $\tilde{{\mathbb{X}}}$ is compact, so $f$ is bounded and attains its bounds almost surely. It turns out that the objective $\overline{f}$ also attains its bounds, so our optimization problem is well-posed.
@@
The objective $\overline{f}$ is bounded and attains its bounds almost surely. Furthermore, there exists a choice ${\mathbf{x}}\in{\mathbb{X}},{\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}$ maximizing (resp. minimizing) $\overline{f}$ such that ${\mathbf{g}}$ has Borel-measurable sample paths.
@@
By construction, the pair ${\mathbf{x}}^{*},{\mathbf{g}}^{*}$ are an upper bound for $\overline{f}$.
@@
In order to argue rigorously about the knowledge gradient, we must make the formula in Equation 9 concrete. Let ${\textnormal{v}}_{\tilde{{\bm{x}}}}=f(\tilde{{\bm{x}}})+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ represent a possible noisy observation at $\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}$. For each $n$, let $\overline{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,\tilde{{\bm{x}}},{\textnormal{v}}_{\tilde{x}})=\mathbb{E}[\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\,|\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}]$ be the expectation of $\overline{f}$ conditional on the $n$ observations so far, and the proposed new observation ${\textnormal{v}}_{\tilde{{\bm{x}}}}$ taken at $\tilde{{\bm{x}}}$. Note that $\overline{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime})=\overline{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,\tilde{{\mathbf{x}}}^{n+1},{\textnormal{v}}^{n+1})$, so this is consistent with our notation from the main text. With this notation, Equation 9 becomes
$\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\mathbb{E}\Biggl{[}\,\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\overline{\mu}^{n+1}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime};\,\tilde{{\bm{x}}},{\textnormal{v}}_{\tilde{{\bm{x}}}})\,\Bigg{|}\,\mathcal{F}^{n}\Biggr{]}-\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\overline{\mu}^{n}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}).$ (25)
We are justified in using a maximum instead of a supremum here because the posterior means $\overline{\mu}^{n}$ are all bounded and attain their bounds.
@@
For all $n$, the posterior mean $\overline{\mu}^{n}$ is bounded and attains its bounds. Furthermore, there exists a choice ${\mathbf{x}}^{*n}\in{\mathbb{X}}$ and ${\mathbf{g}}^{*n}:{\mathbb{U}}\to{\mathbb{Y}}$ maximizing (resp. minimizing) $\overline{\mu}^{\infty}$ such that ${\mathbf{g}}^{*n}$ has Borel-measurable sample paths.
@@
This result also means that our recommendations ${\mathbf{x}}^{*n},{\mathbf{g}}^{*n}\in\operatorname*{arg\,max}_{{\mathbf{x}},{\mathbf{g}}}\overline{\mu}^{n}({\mathbf{x}},{\mathbf{g}})$ from Equation 8 in the main text are well defined.
@@
We conclude this section with a short proof that the joint knowledge gradient is non-negative. This was stated as Lemma 3.1 in the main text.
Proposition A.4 (Lemma 3.1 from the main text).
@@
$\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}\geq\mathbb{E}\bigl{[}\overline{f}({\bm{x}}^{\prime\prime},{\bm{g}}^{\prime\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}.$
@@
$\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}})\bigr{]}\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}\geq\mathbb{E}\bigl{[}\overline{f}({\bm{x}}^{\prime\prime},{\bm{g}}^{\prime\prime})\,\big{|}\,\mathcal{F}^{n}\bigr{]}=\overline{\mu}^{n}({\bm{x}}^{\prime\prime},{\bm{g}}^{\prime\prime}).$
@@
$\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}\geq\max_{{\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime}}\overline{\mu}^{n}({\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime})$
$\Rightarrow\qquad\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}-\max_{{\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime}}\overline{\mu}^{n}({\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime})\geq 0.$
@@ A.2 Convergence of $\alpha_{\text{jKG}}^{n}$ to zero
For $n\in\mathbb{N}_{0}$, the residual uncertainty associated with the joint knowledge gradient is the expected difference between the maximum of the conditional expectation of $\overline{f}$ and the maximum of $\overline{f}$ itself,
$H^{n}=\mathbb{E}\Biggl{[}\,\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\Bigg{|}\,\mathcal{F}^{n}\Biggr{]}-\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\underbrace{\mathbb{E}\bigl{[}\,\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n}\bigr{]}}_{\overline{\mu}^{n}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})}.$ (27)
It measures how well the maximum of the posterior mean $\overline{\mu}^{n}$ approximates the maximum of $\overline{f}$.
@@
Indeed, the first term, $\mathbb{E}[\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,|\,\mathcal{F}^{n}]$, cancels and we have
$H^{n}-\mathbb{E}[H^{n+1}\,|\,\mathcal{F}^{n}]=\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\overline{\mu}^{n+1}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}-\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\overline{\mu}^{n}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})=\alpha_{\text{jKG}}^{n}(\tilde{{\mathbf{x}}}^{n+1}).$
@@
Let $n\in\mathbb{N}_{0}$. We will first show that $|H^{n}|<\infty$ almost surely. We will bound each term separately by the expected maximum of the absolute value of $f$, then show that this is finite. Recalling that $\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})=\mathbb{E}_{{\mathbf{u}}^{\prime}}[f({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\mathbf{u}}^{\prime})]$, we have for the first term in Equation 27,
$\biggl{|}\mathbb{E}\Bigl{[}\,\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}\biggr{|}$ $\leq\mathbb{E}\Bigl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bigl{|}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\bigr{|}\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}$
@@
$\biggl{|}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n}\bigr{]}\biggr{|}$ $\leq\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\Bigl{[}\bigl{|}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\bigr{|}\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}$
@@
$|H^{n}|\leq\biggl{|}\mathbb{E}\Bigl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}\biggr{|}+\biggl{|}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}[\overline{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,|\,\mathcal{F}^{n}]\biggr{|}\leq 2\,\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime}}|f({\mathbf{x}}^{\prime},{\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime})|\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}.$

### v2

@@ A.1 Notation, statistical model and assumptions
$\bar{f}({\bm{x}},{\bm{g}})=\mathbb{E}_{\mathbf{u}}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})].$ (24)
@@
It will be useful later when talking about the posterior mean of $\bar{f}$ to be able to freely interchange the order of integration over $f$ and ${\mathbf{u}}$. This is possible by Fubini’s Theorem / the law of total expectation, provided that the expectation $\mathbb{E}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})]$ over both jointly is well defined.
@@
We can exchange the order of integration over $f$ and ${\mathbf{u}}$ when calculating the posterior mean of $\bar{f}$. That is,
$\bar{\mu}^{n}({\bm{x}},{\bm{g}})=\mathbb{E}\bigl{[}\,\bar{f}({\bm{x}},{\bm{g}})\,\big{|}\,\mathcal{F}^{n}\bigr{]}=\mathbb{E}_{\mathbf{u}}\bigl{[}\mu^{n}({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})\bigr{]}.$
@@
The sample paths of $f$ are continuous and $\tilde{{\mathbb{X}}}$ is compact, so $f$ is bounded and attains its bounds almost surely. It turns out that the objective $\bar{f}$ also attains its bounds, so our optimization problem is well-posed.
@@
The objective $\bar{f}$ is bounded and attains its bounds almost surely. Furthermore, there exists a choice ${\mathbf{x}}\in{\mathbb{X}},{\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}$ maximizing (resp. minimizing) $\bar{f}$ such that ${\mathbf{g}}$ has Borel-measurable sample paths.
@@
By construction, the pair ${\mathbf{x}}^{*},{\mathbf{g}}^{*}$ are an upper bound for $\bar{f}$.
@@
In order to argue rigorously about the knowledge gradient, we must make the formula in Equation 9 concrete. Let ${\textnormal{v}}_{\tilde{{\bm{x}}}}=f(\tilde{{\bm{x}}})+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ represent a possible noisy observation at $\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}$. For each $n$, let $\bar{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,\tilde{{\bm{x}}},{\textnormal{v}}_{\tilde{x}})=\mathbb{E}[\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\,|\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}]$ be the expectation of $\bar{f}$ conditional on the $n$ observations so far, and the proposed new observation ${\textnormal{v}}_{\tilde{{\bm{x}}}}$ taken at $\tilde{{\bm{x}}}$. Note that $\bar{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime})=\bar{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,\tilde{{\mathbf{x}}}^{n+1},{\textnormal{v}}^{n+1})$, so this is consistent with our notation from the main text. With this notation, Equation 9 becomes
$\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\mathbb{E}\Biggl{[}\,\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bar{\mu}^{n+1}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime};\,\tilde{{\bm{x}}},{\textnormal{v}}_{\tilde{{\bm{x}}}})\,\Bigg{|}\,\mathcal{F}^{n}\Biggr{]}-\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bar{\mu}^{n}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}).$ (25)
We are justified in using a maximum instead of a supremum here because the posterior means $\bar{\mu}^{n}$ are all bounded and attain their bounds.
@@
For all $n$, the posterior mean $\bar{\mu}^{n}$ is bounded and attains its bounds. Furthermore, there exists a choice ${\mathbf{x}}^{*n}\in{\mathbb{X}}$ and ${\mathbf{g}}^{*n}:{\mathbb{U}}\to{\mathbb{Y}}$ maximizing (resp. minimizing) $\bar{\mu}^{\infty}$ such that ${\mathbf{g}}^{*n}$ has Borel-measurable sample paths.
@@
This result also means that our recommendations ${\mathbf{x}}^{*n},{\mathbf{g}}^{*n}\in\operatorname*{arg\,max}_{{\mathbf{x}},{\mathbf{g}}}\bar{\mu}^{n}({\mathbf{x}},{\mathbf{g}})$ from Equation 8 in the main text are well defined.
@@
We conclude this section with a short proof that the joint knowledge gradient is non-negative. This was stated as Proposition 3.1 in the main text.
Proposition A.4 (Proposition 3.1 from the main text).
@@
$\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}\geq\mathbb{E}\bigl{[}\bar{f}({\bm{x}}^{\prime\prime},{\bm{g}}^{\prime\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}.$
@@
$\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}})\bigr{]}\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}\geq\mathbb{E}\bigl{[}\bar{f}({\bm{x}}^{\prime\prime},{\bm{g}}^{\prime\prime})\,\big{|}\,\mathcal{F}^{n}\bigr{]}=\bar{\mu}^{n}({\bm{x}}^{\prime\prime},{\bm{g}}^{\prime\prime}).$
@@
$\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}\geq\max_{{\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime}}\bar{\mu}^{n}({\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime})$
$\Rightarrow\qquad\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\bigr{]}\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}-\max_{{\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime}}\bar{\mu}^{n}({\mathbf{x}}^{\prime\prime},{\mathbf{g}}^{\prime\prime})\geq 0.$
@@ A.2 Convergence of $\alpha_{\text{jKG}}^{n}$ to zero
For $n\in\mathbb{N}_{0}$, the residual uncertainty associated with the joint knowledge gradient is the expected difference between the maximum of the conditional expectation of $\bar{f}$ and the maximum of $\bar{f}$ itself,
$H^{n}=\mathbb{E}\Biggl{[}\,\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\Bigg{|}\,\mathcal{F}^{n}\Biggr{]}-\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\underbrace{\mathbb{E}\bigl{[}\,\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n}\bigr{]}}_{\bar{\mu}^{n}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})}.$ (27)
It measures how well the maximum of the posterior mean $\bar{\mu}^{n}$ approximates the maximum of $\bar{f}$.
@@
Indeed, the first term, $\mathbb{E}[\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,|\,\mathcal{F}^{n}]$, cancels and we have
$H^{n}-\mathbb{E}[H^{n+1}\,|\,\mathcal{F}^{n}]=\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bar{\mu}^{n+1}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}-\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bar{\mu}^{n}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})=\alpha_{\text{jKG}}^{n}(\tilde{{\mathbf{x}}}^{n+1}).$
@@
Let $n\in\mathbb{N}_{0}$. We will first show that $|H^{n}|<\infty$ almost surely. We will bound each term separately by the expected maximum of the absolute value of $f$, then show that this is finite. Recalling that $\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})=\mathbb{E}_{{\mathbf{u}}^{\prime}}[f({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\mathbf{u}}^{\prime})]$, we have for the first term in Equation 27,
$\biggl{|}\mathbb{E}\Bigl{[}\,\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}\biggr{|}$ $\leq\mathbb{E}\Bigl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bigl{|}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\bigr{|}\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}$
@@
$\biggl{|}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\bigl{[}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\big{|}\,\mathcal{F}^{n}\bigr{]}\biggr{|}$ $\leq\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}\Bigl{[}\bigl{|}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\bigr{|}\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}$
@@
$|H^{n}|\leq\biggl{|}\mathbb{E}\Bigl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,\Big{|}\,\mathcal{F}^{n}\Bigr{]}\biggr{|}+\biggl{|}\max_{{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime}}\mathbb{E}[\bar{f}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\,|\,\mathcal{F}^{n}]\biggr{|}\leq 2\,\mathbb{E}\biggl{[}\max_{{\mathbf{x}}^{\prime},{\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime}}|f({\mathbf{x}}^{\prime},{\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime})|\,\bigg{|}\,\mathcal{F}^{n}\biggr{]}.$